OpenStudy (anonymous):

Please help I can kind of do this but, then I get stuck!! Ugh!! 2x- 8y = 32 slope-intercept form: slope = y-intercept =

OpenStudy (anonymous):

Slope = 1/4

OpenStudy (anonymous):

y-intercept = -4

OpenStudy (anonymous):

where did you get -4? and 1/4?

OpenStudy (anonymous):

https://www.desmos.com/calculator

OpenStudy (anonymous):

Slope-Intercept Form = y = 1/4x - 4

OpenStudy (texaschic101):

slope intercept form is y = mx + b 2x - 8y = 32 -- subtract 2x from both sides -8y = -2x + 32 -- divide by -8 y = 1/4x - 4 ===> slope intercept form in y = mx + b form, the slope is in the m position, therefore, the slope is 1/4 <==slope in y = mx + b form, the y intercept is in the b position, therefore, the y intercept is -4, or (0,-4) to find the x intercept, sub in 0 for y 2x - 8y = 32 2x - 8(0) = 32 2x = 32 x = 16 <== x intercept is 16 or (16,0)

OpenStudy (texaschic101):

another way to find the y intercept is to sub in 0 for x

OpenStudy (anonymous):

alright give me a moment to try and work this out for myself ok texachic101?

OpenStudy (texaschic101):

ok..sorry

OpenStudy (anonymous):

Its ok thanks you explain well! :)

OpenStudy (texaschic101):

you want me to give you an example question for you to answer ?

OpenStudy (anonymous):

ya hold on let me try and figure out what you put above first because at the beginning I am having trouble finding where you subtracted 2x from both sides because I don't see it.

OpenStudy (texaschic101):

let me put in ALL the steps I used.. 2x - 8y = 32 -- we want this in y = mx + b form, so we have to subtract 2x from both sides 2x - 2x - 8y = -2x + 32 -- combine like terms -8y = -2x + 32 (see how the 2x's cancelled each other out on the right side) now divide both sides by -8, and remember what you do on one side of the equal sign, you have to do on the other side -8y/-8 = -2/-8 x + 32/-8 y = 1/4x - 4 does that help ?

OpenStudy (anonymous):

ya that does help now I am almost done doing it the way you did so that I can understand

OpenStudy (texaschic101):

now when you have it in y = mx + b form, the slope is in the m position and the y intercept is in the b position. slope = 1/4 y intercept = -4 you can also find the y intercept by subbing in 0 for x in the original equation to find the x intercept, sub in 0 for y in the original equation

OpenStudy (anonymous):

Ok I get what you did perfectly!! :)

OpenStudy (anonymous):

It would help if you could post a practice question like this one to help me understand better but, before you do I was going to ask you what the slope intercept was if I already have the slope and y intercept filled out.

OpenStudy (texaschic101):

slope intercept form is y = mx + b if you only needed to find the x and y intercepts, you would not have to put it in y = mx + b form, you would just do your subbing in 0 for each variable to find them. 2x + 4y = 8 turn into slope intercept formula slope = y int = x int = now what can you do with this ?

OpenStudy (anonymous):

ok let me try to figure this out and I will post my answers in a few

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

Do I subtract 4y from both sides like we did before?

OpenStudy (texaschic101):

no...you want to get the y term on one side and everything else on the other side. So by subtracting 2x from both sides, we will end up with 4y on one side and everything else on the other side. So, try subtracting 2x from both sides and see what you get ? However, you can solve it by subtracting 4y from both sides, but then you would have to also subtract 8 from both sides to get the y term by itself

OpenStudy (anonymous):

ok can you like draw this so that I can see it because I don't really know which two sides I should subtract from.

OpenStudy (texaschic101):

what you do on one side of the equal sign, you have to do on the other side. 2x + 4y = 8 -- subtract 2x from both sides 2x - 2x + 4y = -2x + 8 -- combine like terms 4y = -2x + 8 with me so far ?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

Now we divide both sides by 4?

OpenStudy (texaschic101):

here's a little tip....when adding or subtracting a number....when you move it across the equal sign, the sign changes. 3x + 6y = 9....if you wanted to move the 3x across the equal sign, just move it and change the sign. 6y = -3x + 9 or -3x + 6y = 9....if you wanted the -3x moved across the equal sign, move it and change the sign 6y = 3x + 9 this does not work for division/multiplication

OpenStudy (texaschic101):

yes...divide both sides by 4

OpenStudy (anonymous):

ok now I have y=-2/4x+2 correct?

OpenStudy (texaschic101):

correct..but it can be reduced....the slope can

OpenStudy (anonymous):

If this is correct then the y intercept is 2 and the slope is 2 as well if it is reduced and the y=-2/4x+2 is the slope intercept form.

OpenStudy (anonymous):

this is right, right?

OpenStudy (texaschic101):

y = -2/4x + 2 -- reduce any numbers that can be reduced y = -1/2x + 2 y = mx + b -- slope is in m position, y int is in b position y = -1/2x + 2 slope = -1/2 y int = 2 do you see now ?

OpenStudy (anonymous):

oh I get it thanx!! :) Could you help me with other ones I am not 100% ready yet lol

OpenStudy (texaschic101):

sure

OpenStudy (anonymous):

ok hold on I need to get back on my page because I lost it

OpenStudy (texaschic101):

just post it when your ready

OpenStudy (anonymous):

3x - y = 14 slope-intercept form: slope = y-intercept =

OpenStudy (anonymous):

I don't know how to do this is there is not number next to the y

OpenStudy (texaschic101):

when there is not a number...it is and implied number. That just means y is the same as 1y or x is the same as 1x........or -y is the same as -1y...get it ?

OpenStudy (texaschic101):

if it helps, put the number in there.. 3x - 1y = 14

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so I subtract y from both sides? the problem is I don't know which side.

OpenStudy (texaschic101):

3x - 1y = 14.....we want -1y on one side and everything else on the other side. so subtract 3x from both sides...or remember what I said earlier about adding/subtracting...if you want 3x moved to the other side, then move it and change the sign 3x - 1y = 14 -1y = -3x + 14 understand ?

OpenStudy (anonymous):

ok let me try that

OpenStudy (anonymous):

do I subtract 3x from 3x and from 1y or 14?

OpenStudy (texaschic101):

subtract 3x from both sides 3x - 1y = 14 3x - 3x - 1y = -3x + 14 -- combine like terms -1y = -3x + 14 -- the 3x's on the left side cancel out

OpenStudy (anonymous):

so now what?

OpenStudy (anonymous):

do we divide?

OpenStudy (texaschic101):

now we have : -1y = -3x + 14 -- divide both sides by -1

OpenStudy (anonymous):

Is this correct? : y=-3/-1x+-14?

OpenStudy (texaschic101):

correct.....but remember that a negative divided by a negative equals a positive so it would be : y = 3x - 14

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

so that is the slope intercept form?

OpenStudy (texaschic101):

yep..thats the slope intercept form

OpenStudy (texaschic101):

and in y = mx + b form, the slope is in the m position and the y intercept is in the b position. So can you tell me the slope and y int ?

OpenStudy (texaschic101):

y = mx + b y = 3x - 14 slope(m) = y int (b) =

OpenStudy (anonymous):

slope is 3x? and y int is -14?

OpenStudy (texaschic101):

slope is 3 and y int is -14

OpenStudy (texaschic101):

see..your getting this :)

OpenStudy (anonymous):

ok ya I think I am getting it but, sometimes you can see that I do ask questions and get stuck

OpenStudy (anonymous):

but, with you helping me I think I can improve little by little

OpenStudy (texaschic101):

its good to ask questions...it will help you learn

OpenStudy (texaschic101):

y = 4x + 10 slope(m) = y int (b) =

OpenStudy (anonymous):

thanks! :) I have an idea!

OpenStudy (texaschic101):

speak your mind :)

OpenStudy (anonymous):

I have a collaborative chatroom and I can invite you to it if you want so that we could do it there

OpenStudy (texaschic101):

I have never been in the chat room on OS....how embarrassing I feel

OpenStudy (anonymous):

lol it is a scribblar account I got it off of my school thingy I think it's called web2.0 tools

OpenStudy (anonymous):

you want the link?

OpenStudy (anonymous):

and its not on openstudy

OpenStudy (texaschic101):

crap...I am so sorry....I lost track of time...I have to go now. I will be back on later. I am soooo sorry give me the link...please

OpenStudy (anonymous):

ok http://www.scribblar.com/4321rx7t \ I hope you come back soon! :)

OpenStudy (texaschic101):

I will....I always do...lol thanks for the link

OpenStudy (anonymous):

when you are ready just pop in and I will hear a click when you do

OpenStudy (anonymous):

Thanks for all your help and talk to you soon! :)

OpenStudy (texaschic101):

cool....sounds good....I will probably show up there a little later :)

OpenStudy (anonymous):

ok Bye!

OpenStudy (anonymous):

I don't know when you may get there but, if I am not there then I am sorry it must be too late I don't know how long I will wait but, today my teacher is coming

OpenStudy (anonymous):

I hope you do well! :)

OpenStudy (anonymous):

Can you help me with some I have right now while my teacher is not coming til 1 pm

OpenStudy (anonymous):

Me?

OpenStudy (anonymous):

ya I am working on slope intercept and it would really help if I had some extra help

OpenStudy (anonymous):

It depends on what you need help with.

OpenStudy (anonymous):

I probably can help with that

OpenStudy (lstrasmann):

Slope intercept form is y = mx + b where m is the slope and b is the y-intercept. So in your equation, you have to solve for y.

OpenStudy (anonymous):

-y=3x

OpenStudy (anonymous):

slope-intercept form: slope = y-intercept =

OpenStudy (anonymous):

^ the formula. It's correct

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Never mind. It posted slow. http://www.mathwarehouse.com/algebra/linear_equation/slope-of-a-line.php

OpenStudy (anonymous):

ok I think I know enough to do this on my own so I am going to close the question. Thank you anyway! :)

OpenStudy (anonymous):

Your welcome! I hope the best for you! :)

OpenStudy (anonymous):

Thanks! :)

OpenStudy (lstrasmann):

2x-8y=32 -2x -2x -8y=32-2x

OpenStudy (anonymous):

I'm your "fan" now :)